Write the first five terms of the follownig sequence.
a sub n = n + 1/ n with the condition: a sub n= a sub n-1 + 2, where a sub n= 1

Answer by venugopalramana(3286)

Write the first five terms of the follownig sequence.
a sub n = n + 1/ n IS AN INDEPENDENT PROBLEM..FOR THIS
A1=1+1/1=2
A2=2+1/2=5/2
A3=3+1/3=10/3
A4=4+1/4=17/4
A5=5+1/5=26/5
with the condition: a sub n= a sub n-1 + 2, where a sub n= 1....THIS IS A SEPERATE PROBLEM....
A1=1
A2=A1+2=1+2=3
A3=A2+2=3+2=5
A4=A3+2=5+2=7
A5=A4+2=7+2=9